OpenStudy (anonymous):

evaluate 3 + 4 (1/2-3/4)

OpenStudy (anonymous):

\[3+4(\frac{ 1 }{ 2 }-\frac{ 3 }{ 4 })=3+4(\frac{ 2 }{ 4 }-\frac{ 3 }{ 4 })=3+4(\frac{ -1 }{ 4 })=3+(-1)=3-1=2\]

OpenStudy (anonymous):

Thank you mononoor